# On the best approximation by finite Gaussian mixtures

### v1

@@ 3.1 Compactly supported distributions
Denote $\mu=\mathbb{E}[Q]$ and $\sigma^{2}=\mathsf{var}[Q]$. Let $U\sim P$, $V\sim Q$, $U-\mu\sim{P}^{\prime}_{\mu}$, and $V-\mu\sim{Q}^{\prime}_{\mu}$. Since $m_{j}(P)=m_{j}(Q)$ for $j\in[J-1]$, it holds that $m_{j}({P}^{\prime}_{\mu})=m_{j}({Q}^{\prime}_{\mu})$ for $j\in[J-1]$. Applying [WY20, Lemma 9] with $\mathbb{E}_{{Q}^{\prime}_{\mu}}[X]=0$, we have
@@
where the last inequality follows from the Poisson tail bound $\mathbb{P}[X\geq J]\leq e^{-4M^{2}}\left(\frac{4eM^{2}}{J}\right)^{J}$ for $X\sim\operatorname{Poisson}(4M^{2})$ and $J>4M^{2}$ [Ver18, Exercise 2.3.3]. ∎
Now we are ready to show the $\chi^{2}$ approximation error bound for compactly supported mixing distributions. When $M$ is small, we construct a discrete distribution that matches the first few moments. However, this approach is loose for large $M$ and can be remedied using the idea of local approximation. Specifically, we partition the support into subintervals and approximate each conditional distribution via moment matching; similar construction of local moment matching has appeared in the statistics literature [Zha09, SW23] by applying Caratheódory’s theorem to the truncated moments. The overall approximation error can be bounded combining that in each subinterval.
@@ 3.2 Distribution families under tail conditions
In this subsection, we extend our analysis to the sub-Weibull family ${\mathcal{P}}_{\alpha}(\beta)$ using a truncation argument. For each $P\in{\mathcal{P}}_{\alpha}(\beta)$, we reduce the problem to approximating the conditional distribution $P_{t}$ of $P$ on $[-t,t]$, where the approximate error is bounded by Theorem 3 below. The final result follows by optimizing $t$. Moreover, this approach is applicable for general distribution families by incorporating the corresponding tail probability bounds.
@@
${\mathcal{E}}^{\star}(m,{\mathcal{P}}_{2}(\sigma),\chi^{2})\leq\exp\left(-Cm\log\left(1+\frac{1}{\sigma}\right)\right).$
@@ 4 Converse via spectrum of moment matrices
In this section, we give lower bounds of the best approximation error by finite mixtures. In Section 4.1, we first propose a general framework applicable to any finite mixture in terms of trigonometric moment matrices. This strategy is then applied to the class of sub-Weibull distributions ${\mathcal{P}}_{\alpha}(\beta)$ and compactly supported distributions ${\mathcal{P}}^{\textsf{Bdd}}_{M}$ in Sections 4.2 and 4.3, respectively. For each distribution family ${\mathcal{P}}$ of interest, we exhibit a test distribution $P\in{\mathcal{P}}$ and obtain lower bound of ${\mathcal{E}}^{\star}(\epsilon,P,d)$ and hence also ${\mathcal{E}}^{\star}(m,{\mathcal{P}},d)$. These test distributions are essentially the hardest to approximate in their respective distribution classes, and are amenable to analysis. Finally, in Section 4.4 we focus on the regime of large $\beta$ and $M$ and prove lower bound on $m$ such that consistent approximation is possible.
@@ 4.1 General framework
We state a general procedure to lower bound the approximation error ${\mathcal{E}}^{\star}(m,P,d)$, where $P$ is not finitely supported. At the high level, we relate the approximation error in TV to that of the trigonometric moments. As mentioned in Section 2.1, the trigonometric moment matrix of $P$ of any degree is full-rank, while that of any $m$-atomic distribution has rank at most $m$. Thus, the question boils down to the smallest eigenvalue of the trigonometric moment matrix of $P$, which can be lower bounded using classical results in spectral analysis. For this step, we provide two different methods, both of which are needed for proving the main results.
@@ 4.2 Distribution families under tail conditions
Denote $X\sim P_{\alpha,\beta}$ and $Y=\delta X\sim g$, where $g(\theta)=\frac{1}{\delta}f_{\alpha,\beta}(\frac{\theta}{\delta})$. By Proposition 2 and (17), we have
${\mathcal{E}}^{\star}(m,P_{\alpha,\beta},\mathsf{TV})\geq\sup_{\delta>0}\frac{\lambda_{\min}({\mathbf{T}}_{m}(Y))}{2(m+1)\exp(m^{2}\delta^{2}/2)}\geq\sup_{\delta>0}\frac{\pi\min_{0\leq\theta\leq 2\pi}g^{\mathsf{wrap}}(\theta)}{(m+1)\exp(m^{2}\delta^{2}/2)}.$ (21)
Note that $g^{\mathsf{wrap}}(\theta)\geq\inf_{\theta\in[0,2\pi)}g(\theta)=\frac{1}{\delta}f_{\alpha,\beta}\left(\frac{2\pi}{\delta}\right)$. Choosing $\delta=\left(\frac{1}{m}\right)^{\frac{2}{2+\alpha}}\left(\frac{2\pi}{\beta}\right)^{\frac{\alpha}{2+\alpha}}$, we have
@@
where the explicit constants are obtained from (21) with $g(\theta)=\frac{1}{\delta}f(\frac{\theta}{\delta})$. [^3] Slightly different from the proof of Theorem 5, here we lower bound the wrapped density with $g^{\mathsf{wrap}}(\theta)\geq\inf_{\theta\in[0,\pi]}g(\theta)$ for symmetric distributions, which improves the multiplicative constants by a factor of 2. Similarly, for the $N(0,\sigma^{2})$ distribution with $(\alpha,\beta)=(2,\sqrt{2}\sigma)$,
@@
4.3 Compactly supported distributions
We focus on the class ${\mathcal{P}}^{\textsf{Bdd}}_{M}$ of all distributions supported on $[-M,M]$. For $M\lesssim m\lesssim M^{2}$, the desired lower bound can be directly obtained by applying Proposition 2 to the trigonometric moment matrix of the uniform distribution. However, the case of $m\gtrsim M^{2}$ is more challenging, which requires a small frequency parameter $\delta$. For $X$ supported on $[-M,M]$, $\delta X\mod 2\pi\mathbb{Z}$ is the same as as $\delta X$ for small $\delta$. As such, the minimum value of the wrapped density function is zero, which only yields a trivial lower bound. Instead, we apply the second approach based on orthogonal expansion to distributions supported on an arc of the unit circle.
Lemma 4.
For any distribution $P$ on $\mathbb{R}$, under the notations above,
$\|{\mathbf{R}}_{m}\|_{F}^{2}\leq\sum_{n=0}^{m}2^{2n}\kappa_{n}^{2}.$
Proof.
It follows that
[…]
$\int p_{j}(x)p_{k}(x)w(x)\mathrm{d}x=h_{j}{\mathbf{1}{\left\{{j=k}\right\}}}.$
Lemma 5.
Now, with Lemmas 4 and 5, we are ready to show the lower bound for ${\mathcal{P}}^{\textsf{Bdd}}_{M}$:
@@ 4.3 Compactly supported distributions
For any $M>0$ and $m\in\mathbb{N}$,
Furthermore, there exists a universal constant $C_{1}$ and $C^{\prime}$ such that when $m\geq C^{\prime}M^{2}$,
${\mathcal{E}}^{\star}(m,{\mathcal{P}}^{\textsf{Bdd}}_{M},\mathsf{TV})\geq\exp\left(-C_{1}m\log\left(\frac{m}{M^{2}}\right)\right).$ (24)
Proof.
First, we prove (23). Let $X\sim P=\mathrm{Unif}[-M,M]$ and $\delta=\frac{\pi}{M}$. Then $\delta X\sim\mathrm{Unif}[-\pi,\pi]$, and $t_{k}(\delta X)=\mathbb{E}[e^{ik\delta X}]={\mathbf{1}{\left\{{k=0}\right\}}}$ for $k\in\mathbb{Z}$. Consequently, ${\mathbf{T}}={\mathbf{T}}_{m}(\delta X)=I_{m+1}$ is the identity matrix and so $\lambda_{\min}({\mathbf{T}})=1$. By Proposition 2,
Next, we prove (24) by constructing a distribution supported on $[-M,M]$. Let $b\in(0,\pi)$ and $\gamma=\sin\frac{b}{2}$. Let $\tilde{P}$ be the distribution supported on $[-b,b]$ with a density function
Let $\{\Psi_{n}\}$ and $\{\tilde{\varPhi}_{n}\}$ be the sets of monic orthogonal polynomials on the unit circle associated with $Q$ and $\tilde{P}$ as defined in (12), and $\psi_{n}=\kappa_{n}\Psi_{n}$, $\tilde{\varphi}_{n}=\tilde{\kappa}_{n}\tilde{\varPhi}_{n}$ be the orthonormal polynomials, respectively. By definition, $\tilde{\varPhi}_{n}(z)=(-1)^{n}\Psi_{n}(-z)$ and $\tilde{\kappa}_{n}=\kappa_{n}$. Let $\mathbf{\tilde{R}}_{m}$ be the coefficient matrix of $\{\tilde{\varPhi}_{n}\}$ of order $m+1$. By Lemmas 4 and 5, we obtain that
$r_{n}(x)=x-\frac{1}{4r_{n-1}(x)},\quad n\in\mathbb{N}.$
$\|\tilde{{\mathbf{R}}}_{m}\|_{F}^{2}\leq\sum_{n=0}^{m}2^{2n+1}\gamma^{-2n}=2\frac{(2/\gamma)^{2m+2}-1}{(2/\gamma)^{2}-1}.$
We conclude the proof of (24) by plugging in the choice of $b$. ∎
Remark 3.
Although (24) is proved by choosing a non-uniform distribution on $[-M,M]$, we can show that this lower bound also holds for $\mathrm{Unif}[-M,M]$. For this result, we first lower bound the approximation error in $\chi^{2}$, which can be related to, via the variational representation of the $\chi^{2}$-divergence, to the smallest eigenvalue of a certain weighted moment matrix. Then, we convert the $\chi^{2}$-divergence lower bound to a $\mathsf{TV}$ lower bound using a truncation argument for Gaussian mixtures. This result in presented in Appendix B.2 – see Theorem 11 therein.
4.4 Inconsistency regime
In this subsection, we consider the regime when the scale parameter $M$ or $\beta$ diverges. In this case, more components are needed to achieve a prescribed approximation accuracy. Hence, it is natural to conjecture that consistent approximation is impossible unless $m$ is at least proportional to the scale parameter. This however is not captured by Theorems 5 and 6 where the lower bound still converges to zero when $m\asymp M$ (or $m\asymp\beta$) due to the polynomial term in $m$.
The next result makes the intuition precise in a strong sense for the test distributions used in the previous subsections. Instead of applying Proposition 2 and bounding the minimum eigenvalue, we execute a different but more straightforward proof strategy: If the mixing distribution $P$ is smooth enough with a large $M$ or $\beta$, then $f_{P}$ can be flat over a large region. In comparison, when $m$ is too small, $f_{P_{m}}$ inevitably has many modes and thus cannot approximate $f_{P}$ consistently.
Theorem 7.
The following statements hold:
  1. 1.
@@ 4.4 Inconsistency regime
$\overset{\rm(a)}{\geq}1-\sum_{i\not\in{\mathcal{I}}}w_{i}-\sum_{i\in{\mathcal{I}}}\frac{4C_{\alpha}}{\beta}\eta(w_{i})$
$\overset{\rm(b)}{\geq}1-m\frac{2\tilde{C}_{\alpha}}{\beta}-\frac{4C_{\alpha}}{\beta}|{\mathcal{I}}|\eta(|{\mathcal{I}}|^{-1})$
@@
$\overset{\rm(c)}{\geq}1-\frac{2\tilde{C}_{\alpha}m}{\beta}-\frac{4}{\sqrt{\pi}}\frac{\tilde{C}_{\alpha}m}{\beta}\sqrt{\log\frac{\beta}{\tilde{C}_{\alpha}m}},$
@@
The lower bound holds uniformly for all $P_{m}\in{\mathcal{P}}_{m}$, and hence the result (25) follows.
For $P=\mathrm{Unif}[-M,M]$, the derivation is essentially the same as above. Note that $f_{P}(x)=\frac{1}{2M}\int_{-M}^{M}\phi(x-\theta)\mathrm{d}\theta\leq\frac{1}{2M}.$ For any $P_{m}=\sum_{i=1}^{m}w_{i}\delta_{\theta_{i}}\in{\mathcal{P}}_{m}$, define $R_{i}=\{x:w_{i}\phi(x-\theta_{i})\geq(2M)^{-1}\}$ and ${\mathcal{I}}\triangleq\{i\in[m]:w_{i}\geq\sqrt{2\pi}M^{-1}\}$ correspondingly. Finally, (26) is proved via a similar argument. ∎

### v2

@@ 3.1 Compactly supported distributions
Denote $\mu=\mathbb{E}[Q]$ and $\sigma^{2}=\mathsf{var}[Q]$. Let $U\sim P$, $V\sim Q$, $U-\mu\sim{P}^{\prime}_{\mu}$, and $V-\mu\sim{Q}^{\prime}_{\mu}$. Since $m_{j}(P)=m_{j}(Q)$ for $j\in[J-1]$, it holds that $m_{j}({P}^{\prime}_{\mu})=m_{j}({Q}^{\prime}_{\mu})$ for $j\in[J-1]$. Applying [WY20a, Lemma 9] with $\mathbb{E}_{{Q}^{\prime}_{\mu}}[X]=0$, we have
@@
where the last inequality follows from the Chernoff bound for Poisson distribution $\mathbb{P}[X\geq J]\leq e^{-4M^{2}}\left(\frac{4eM^{2}}{J}\right)^{J}$ for $X\sim\operatorname{Poisson}(4M^{2})$ and $J>4M^{2}$ (see, e.g., [MU05, Theorem 4.4]). ∎
Now we are ready to show the $\chi^{2}$ approximation error bound for compactly supported mixing distributions. When $M$ is small, we construct a discrete distribution that matches the first few moments. However, this approach is loose for large $M$ and can be remedied using the idea of local approximation. Specifically, we partition the support into subintervals and approximate each conditional distribution via moment matching; similar construction of local moment matching has appeared in the statistics literature [Zha09, SW24] by applying Caratheódory’s theorem to the truncated moments. This approach determines the allocation of atoms across subintervals, and tight upper bounds then follow from the trade-off between the number of atoms per interval and the interval size. The overall approximation error can be bounded combining that in each subinterval.
@@ 3.2 Distribution families under tail conditions
In this subsection, we extend our analysis to the sub-Weibull family ${\mathcal{P}}_{\alpha}(\beta)$ using a truncation argument. For each $P\in{\mathcal{P}}_{\alpha}(\beta)$, we reduce the problem to approximating the conditional distribution $P_{t}$ of $P$ on $[-t,t]$, where the approximate error is bounded by Theorem 3 above. The final result follows by optimizing $t$. Moreover, this approach is applicable for general distribution families by incorporating the corresponding tail probability bounds.
@@
${\mathcal{E}}^{\star}(m,{\mathcal{P}}_{2}(\sigma),\chi^{2})\leq\exp\left(-Cm\log\left(1+\frac{1}{\sigma}\right)\right).$ (17)
@@ 4 Converse via spectrum of moment matrices
In this section, we give lower bounds of the best approximation error by finite mixtures. In Section 4.1, we first propose a general framework applicable to any finite mixture in terms of trigonometric moment matrices. This strategy is then applied to the class of sub-Weibull distributions ${\mathcal{P}}_{\alpha}(\beta)$ and compactly supported distributions ${\mathcal{P}}^{\textsf{Bdd}}_{M}$ in Sections 4.2 and 4.3, respectively. For each distribution family ${\mathcal{P}}$ of interest, we exhibit a test distribution $P\in{\mathcal{P}}$ and obtain lower bound of ${\mathcal{E}}^{\star}(\epsilon,P,d)$ and hence also ${\mathcal{E}}^{\star}(m,{\mathcal{P}},d)$. These test distributions are essentially the hardest to approximate in their respective distribution classes and are amenable to analysis.
Our results indicate that distributions with heavier tails are generally more difficult to approximate using finite mixtures. Specifically, the tail behavior itself directly yields a lower bound on the approximation error (see Corollary 1 for an explicit result). In addition, without imposing constraints on the distribution family, no finite mixture of fixed order can achieve consistent approximations. To identify the fundamental limit, we consider the regime of large $\beta$ or $M$ and prove lower bound on $m$ such that consistent approximation is possible (see Theorem 6 and (28) of Theorem 7).
@@ 4.1 General framework
We state a general procedure to lower bound the approximation error ${\mathcal{E}}^{\star}(m,P,d)$, where $P$ is not finitely supported. At the high level, we relate the approximation error in $\mathsf{TV}$ to that of the trigonometric moments. As mentioned in Section 2.1, the trigonometric moment matrix of $P$ of any degree is full-rank, while that of any $m$-atomic distribution has rank at most $m$. Thus, the question boils down to the smallest eigenvalue of the trigonometric moment matrix of $P$, which can be lower bounded using classical results in spectral analysis. For this step, we provide two different methods, both of which are needed for proving the main results.
@@ 4.1.1 Lower bound via trigonometric moments
Remark 3.
Proposition 2 is proved via lower bounding $\mathsf{TV}(f_{P},f_{P_{m}})$ by the difference of trigonometric moments. A related result [DWYZ23, Theorem 4.2] lower bounds the divergence by moment differences as $H^{2}(f_{P},f_{Q})\geq C_{k}\|{\mathbf{m}}_{2k-1}(P)-{\mathbf{m}}_{2k-1}(Q)\|_{2}^{2}$ for any $P,Q\in{\mathcal{P}}^{\textsf{Bdd}}_{1}\cap{\mathcal{P}}_{k}$, where ${\mathbf{m}}_{2k-1}(P)\triangleq(m_{1}(P),\dots,m_{2k-1}(P))$ and $C_{k}=\exp(-\Theta(k\log k))$. However, this result require both $f_{P}$ and $f_{Q}$ to be finite mixtures. In contrast, the lower bound on $\mathsf{TV}$ in Proposition 2 applies to continuous mixture $f_{P}$.
@@ 4.1.2 Spectral analysis of trigonometric moment matrices
As a corollary of Proposition 2, we provide an explicit lower bound based on the tail behavior of the mixing distribution. This result highlights that distributions with heavier tails are more difficult to approximate. We will formalize this intuition in the following subsections.
Corollary 1.
Let $X\sim P$ with density function $f$ on $\mathbb{R}$. Then,
${\mathcal{E}}^{\star}(m,P,\mathsf{TV})\geq\sup_{\delta>0}\frac{\pi\inf_{0\leq\theta\leq 2\pi/\delta}f(\theta)}{(m+1)\delta\exp(m^{2}\delta^{2}/2)}.$ (21)
Moreover, if $f$ is symmetric around zero, then the infimum in the numerator may be take over $0\leq\theta\leq\pi/\delta$.
Proof.
Denote $Y=\delta X\sim g$ with density function $g(\theta)=\frac{1}{\delta}f(\frac{\theta}{\delta})$. Applying (18) with the fact that $g^{\mathsf{wrap}}(\theta)\geq\inf_{\theta\in[0,2\pi)}g(\theta)$, we have $\lambda_{\min}({\mathbf{T}}_{m}(Y))\geq 2\pi\inf_{\theta\in[0,2\pi)}g(\theta)$. Particularly, for symmetric $f$, $\lambda_{\min}({\mathbf{T}}_{m}(Y))\geq 2\pi\inf_{\theta\in[0,\pi)}g(\theta)$. Applying Proposition 2, (21) then follows. ∎
@@ 4.2 Distribution families under tail conditions
Denote $X\sim P_{\alpha,\beta}$ and $Y=\delta X\sim g$, where $g(\theta)=\frac{1}{\delta}f_{\alpha,\beta}(\frac{\theta}{\delta})$. Note that $f_{\alpha,\beta}$ is symmetric and decreasing in $(0,\infty)$. Applying (21) with $\delta=\left(\frac{1}{m}\right)^{\frac{2}{2+\alpha}}\left(\frac{2\pi}{\beta}\right)^{\frac{\alpha}{2+\alpha}}$, we have
@@
where the explicit constants are obtained from (21). Similarly, for the $N(0,\sigma^{2})$ distribution with $(\alpha,\beta)=(2,\sqrt{2}\sigma)$,
@@
Next, we consider the regime when the scale parameter $\beta$ diverges. In this case, more components are needed to achieve a prescribed approximation accuracy. Hence, it is natural to conjecture that consistent approximation is impossible unless $m$ is at least proportional to the scale parameter. This however is not captured by Theorems 5 where the lower bound still converges to zero when $m\asymp\beta$ due to the polynomial term in $m$.
The next result makes the intuition precise in a strong sense for the given test distributions. Instead of applying Proposition 2 and bounding the minimum eigenvalue, we execute a different but more straightforward proof strategy: If the mixing distribution $P$ is smooth enough with a large $\beta$, then $f_{P}$ can be flat over a large region. In comparison, when $m$ is too small, $f_{P_{m}}$ inevitably has many modes and thus cannot approximate $f_{P}$ consistently.
@@
$\overset{\mathrm{(a)}}{\geq}1-\sum_{i\not\in{\mathcal{I}}}w_{i}-\sum_{i\in{\mathcal{I}}}\frac{4C_{\alpha}}{\beta}\eta(w_{i})$
$\overset{\mathrm{(b)}}{\geq}1-m\frac{2\tilde{C}_{\alpha}}{\beta}-\frac{4C_{\alpha}}{\beta}|{\mathcal{I}}|\eta(|{\mathcal{I}}|^{-1})$
@@
$\overset{\mathrm{(c)}}{\geq}1-\frac{2\tilde{C}_{\alpha}m}{\beta}-\frac{4}{\sqrt{\pi}}\frac{\tilde{C}_{\alpha}m}{\beta}\sqrt{\log\frac{\beta}{\tilde{C}_{\alpha}m}},$
@@
The lower bound holds uniformly for all $P_{m}\in{\mathcal{P}}_{m}$, and hence the result (24) follows. ∎
4.3 Compactly supported distributions
We focus on the class ${\mathcal{P}}^{\textsf{Bdd}}_{M}$ of all distributions supported on $[-M,M]$. Theorem 7 shows that the desired lower bound can be obtained by considering the uniform distribution.
Theorem 7.
There exists universal constants $C_{1}$ and $C^{\prime}$ such that for any $m\geq C^{\prime}(M\vee 1)^{2}$,
${\mathcal{E}}^{\star}(m,\mathrm{Unif}[-M,M],\mathsf{TV})\geq\exp\left(-C_{1}m\log\left(\frac{m}{M^{2}}\right)\right),$ (26)
and it holds for any $M>0$, $m\in\mathbb{N}$ that
Moreover, for $C=\sqrt{\frac{\pi}{2}}$ and any $M\geq\sqrt{2\pi}m$,
We give some interpretation of Theorem 7 as follows.
  • Non-vanishing error. When $m\leq o(M)$, (28) gives a compact support analogue of Theorem 6, demonstrating that $m$-component mixtures cannot yield any non-trivial approximation as the TV error is $1-o(1)$. The derivation of (28) is essentially the same as those used in Theorem 6.
  • Polynomially small error. When $M\lesssim m\lesssim M^{2}$, (27) dominates the lower bound. The lower bound can be directly obtained by applying Proposition 2 to the trigonometric moment matrix of the uniform distribution.
  • Exponentially small error. When $m\gtrsim M^{2}$, (26) becomes effective. The proof is more challenging, which requires a small frequency parameter $\delta$ in Proposition 2. For $X$ supported on $[-M,M]$, $\delta X\mod 2\pi\mathbb{Z}$ is the same as $\delta X$ for small $\delta$. As such, the minimum value of the wrapped density function is zero, which only yields a trivial lower bound.
    For this result, we first lower bound the approximation error in $\chi^{2}$, which can be related to, via the variational representation of the $\chi^{2}$-divergence, the smallest eigenvalue of a certain weighted moment matrix. Then, we convert the $\chi^{2}$-divergence lower bound to a $\mathsf{TV}$ lower bound using a truncation argument for Gaussian mixtures.
Proof of Theorem 7.
The proof of (26) is provided in Appendix B.2 – see Theorem 11 therein. Let $X\sim P=\mathrm{Unif}[-M,M]$ and $\delta=\frac{\pi}{M}$. By Corollary 1,
The proof of (28) is essentially the same as the proof of Theorem 6. Note that $f_{P}(x)=\frac{1}{2M}\int_{-M}^{M}\phi(x-\theta)\mathrm{d}\theta\leq\frac{1}{2M}.$ For any $P_{m}=\sum_{i=1}^{m}w_{i}\delta_{\theta_{i}}\in{\mathcal{P}}_{m}$, define
$R_{i}=\{x:w_{i}\phi(x-\theta_{i})\geq(2M)^{-1}\},\quad{\mathcal{I}}\triangleq\{i\in[m]:w_{i}\geq\sqrt{2\pi}M^{-1}\}.$
The result is proved via a similar argument to that of Theorem 6. ∎
For $m\gtrsim M^{2}$, the spectral analysis framework remains applicable if we consider certain compactly supported distribution other than the uniform distribution. This framework will be extended to multiple dimensions in Section 5.4. To illustrate this, let $b\in(0,\pi]$ and $\tilde{P}$ be the distribution supported on $[-b,b]$ with the density function
$\tilde{f}(\theta)=\begin{cases}\frac{\sqrt{\sin^{2}(b/2)-\sin^{2}(\theta/2)}\cos(\theta/2)}{\pi\sin^{2}(b/2)},&\theta\in[-b,b],\\ 0,&\text{otherwise.}\end{cases}$ (29)
Let $X\sim\tilde{P}$ and $\delta=\frac{b}{M}$. Then $\frac{X}{\delta}$ is supported on $[-M,M]$. Let $P$ denote the distribution of $\frac{X}{\delta}$. The following result holds:
Proposition 4.
Let $P$ be defined as above. Then, there exists $b\in(0,\pi]$ and universal constants $C_{1}$, $C^{\prime}$ such that when $m\geq C^{\prime}M^{2}$,
${\mathcal{E}}^{\star}(m,P,\mathsf{TV})\geq\exp\left(-C_{1}m\log\left(\frac{m}{M^{2}}\right)\right).$ (30)
To prove Proposition 4, we connect the orthogonal system on an arc of the unit circle to the orthogonal system on an interval of the real line (see Lemma 9). Specifically, the orthonormal polynomials associated with $\tilde{P}$ can be related to the (scaled) Chebyshev polynomials of the second kind. Propositions 2 and 3 can then be applied, where the explicit form of the Chebyshev polynomials is utilized to bound the norm of the associated coefficient matrix. The complete proof is presented in Appendix B.3.